OpenStudy (anonymous):

find the equation for a parabola with it's focus at (0,3) and a directrix at y = -3

OpenStudy (anonymous):

|a| = 1/(4p) f(x) = a(x - h)^2 + k Plug in the numbers. p is the focal distance. a = 1/(4(3)) a = 1/12 f(x) = 1/12(x - h)^2 + k Since the focus is at (0, 3) and the directrix at y = -3, there is no phase shift or vertical shift, so h and k and disappear. f(x) = 1/12x^2